# Suite-IN++: A FlexiWear BodyNet Integrating Global and Local Motion Features from Apple Suite for Robust Inertial Navigation

### v1

@@ IV-C Weighted global fusion for overall motion trends capture
Inspired by[44], our model adopts a weighted fusion method to estimate the information quality (quality weight) contributed by each device and aggregate the motion features from all wearable devices in a weighted combination manner. Through this operation, our model aims to aggregate the effective global motion information contained in devices from different body parts into a shared low-dimensional latent space, so that the overall motion trend can be more robustly estimated without being affected by flexible and changeable device configurations.
@@ IV-D Attentive local analysis for motion details acquisition
| IONet*[24] | RoNIN*[12] | IMUNet*[41] | DeepIT*[37] | ReWF[36] | Deep Sense[45] | Suite-IN[3] | Suite-IN++ | ||
@@ V Experiment
In this section, we evaluate the proposed method and compare its performance with the state-of-the-art techniques. In all experiments, we keep the dataset settings consistent, with the ratio of training set, validation set, and test set being 6:2:2.
@@ V-A Experimental Setup
The architecture was implemented with PyTorch and trained on a NVIDIA NTX 4070 GPU. We used Adam, a first-order gradient-based optimizer[57], with a learning rate of 0.0001, a batch size of 128, and a window size of 100. On average, the training converged after 100 iterations. To avoid overfitting, we collected data with rich motion features and adopted Dropout [58]in the network, randomly dropping 20% of the units from the neural network during training. We set the hyper-parameters, , , , , , , .
We employ three standard metrics, as proposed in[59], to rigorously evaluate our results:
@@ V-B Compared Algorithms
Pedestrian Dead Reckoning (PDR)[60]: We utilize a step-counting algorithm to detect foot-steps and move the position along the device heading direction by a predefined distance of 0.67m per step.
@@
IONet*[24]: A deep learning-based inertial navigation method employing an LSTM network model. To extend IONet for three-node positioning, we concatenate the data from three devices, referring to this variant as IONet*. Notably, the original IONet regresses distance and heading changes within each window; we adapt this to estimate velocity, ensuring consistency with our method.
RoNIN*[12]: A deep learning-based inertial tracking method that employs three different backbones (LSTM, ResNet, TCN), with ResNet achieving the highest accuracy. We extend the official RoNIN (ResNet) implementation by applying a concatenation approach for three-node data fusion, denoted as RoNIN*.
IMUNet*[41]: IMUNet introduces a one-dimensional version of the state-of-the-art convolutional neural network (CNN) network for inertial position estimation on the edge device implementation. We extend it to three-node positioning through concatenation, named as IMUNet*.
@@
DeepIT*[37]: An inertial navigation method that integrates smartphone and headphone data using an LSTM. We extend it with DeepIT* to fuse data from three sensors with primal weighting and modify the regression target from distance and heading to velocity for consistency with our method.
@@
DeepSense*[45]: DeepSense is the classic learning model for HAR of multi-sensor data. The architecture of DeepSense includes three layers of local CNN, three layers of global CNN and two layers of GRU. We implement DeepSense* for positioning based on the settings in the original article.
Suite-IN[3]: Our previous work, a three-node positioning method based on wearable devices that uses shared global motion information contained in multiple devices.
@@ V-C Comparison with the State-of-the-Art methods
Specifically, in the stable walking mode STW, our method achieves comparable positioning results with the classic single-node localization algorithm IMUNet*[41], and outperforms multiple other localization methods. As our algorithm is specially designed to improve the positioning accuracy for complex walking modes, our method shows significant advantages in challenging scenarios (PVW, MVW, DLW, and DRW). Suite-IN[3]achieves suboptimal performance in overall experiment. Compared with Suite-IN, our method significantly reduces ATE and RTE in all walking modes, ATE and RTE are reduced by 0.12m (3.83%) and 0.44m (11.03%) for STW, 1.02m (31.86%) and 1.33m (31.31%) for PVW, 1.43 m (33.3%) and 1.22 m (25.10%) for MVW, 0.07 m (2.12%) and 0.31 m (8.89%) for DLW, 0.11 m (4.21%) and 0.59 m (15.94%) for DRW. Compared with DeepSense[45], a classic multi-sensor fusion method for human motion analysis, our approach achieves significant improvements even in complex walking modes. For MVW, DLW, and DRW, ATE is reduced by 3.15m, 2.00m, and 3.05m, while RTE decreases by 3.01m, 1.87m, and 3.18m, respectively. These results highlight the effectiveness of our method in localization under complex motion conditions.
@@
| No. | Modules | Data setting | |||||||||||||
| Contrast.FE | Weighted.GF | Attentive.LA | Overall | STW | PVW | MVW | DLW | DRW | |||||||
| ATE | RTE | ATE | RTE | ATE | RTE | ATE | RTE | ATE | RTE | ATE | RTE | ||||
| (1) | 3.229 | 3.627 | 3.279 | 3.435 | 2.397 | 3.548 | 3.717 | 4.133 | 3.689 | 3.828 | 2.436 | 3.184 | |||
| (2) | 3.016 | 3.524 | 3.068 | 3.543 | 2.380 | 3.485 | 3.573 | 4.228 | 3.161 | 3.363 | 2.455 | 2.925 | |||
| (3) | 3.209 | 3.642 | 3.129 | 3.392 | 2.095 | 3.183 | 3.367 | 4.074 | 3.635 | 3.499 | 2.863 | 3.744 | |||
| (4) | 3.383 | 3.740 | 3.352 | 3.633 | 2.387 | 3.222 | 4.113 | 4.124 | 3.569 | 3.477 | 2.978 | 4.085 | |||
| (5) | 4.101 | 4.194 | 3.749 | 3.952 | 3.461 | 4.329 | 5.029 | 4.935 | 5.006 | 4.101 | 3.169 | 3.836 | |||
| (6) | 2.915 | 3.253 | 3.062 | 3.542 | 2.188 | 2.909 | 2.865 | 3.628 | 3.147 | 3.156 | 2.434 | 3.083 | |||
@@ V-D Ablation study
Based on variant 3), 6) introduces the Weighted.GF module to enhance global motion feature utilization. The results show that this module further improves the positioning accuracy in various walking modes, particularly in complex scenarios, as global features offer a more stable motion pattern, mitigating the negative impact of limb shaking on local features. Combining global and local motion information enables a more comprehensive capture of walking characteristics, enhancing robustness in challenging environments.
@@ V-E Tests Involving Multiple Subjects
We selected the single-node positioning method IMUNet*[41] and the multi-node positioning method Suite-IN[3] as benchmark comparisons to evaluate the performance advantages of our method, as they perform outstandingly in their respective categories according to Tab. III. As shown in Fig. 9, our method can still maintain high positioning accuracy on unseen users without significant error growth, while the positioning performance of IMUNet and Suite-IN on unseen users is significantly reduced, indicating that they are less adaptable to the motion patterns of different individuals. In contrast, our method can more comprehensively model and adapt to the motion patterns of different individuals by effectively decoupling and utilizing different motion features, thereby significantly improving the generalization ability of the model.
@@ V-F Comparison of Suite-IN++ and Suite-IN
In this section, we compare two wearable-based pedestrian localization methods: Suite-IN++ and Suite-IN[3]. Unlike our previous work, Suite-IN, which treats local motion features as interference, Suite-IN++ leverages the rich motion information within these features to enhance position estimation.
@@ VI CONCLUSIONS
This paper introduces a flexiwear bodynet-based inertial dataset covering flexible device configurations and complex walking modes, and proposes an innovative inertial positioning method based on real-life wearable devices. By extracting and integrating both global and local motion features, our method effectively captures the overall motion trend and detailed dynamics, leveraging sensor data from various parts of the body to achieve robust and high-precision pedestrian positioning. Experimental results demonstrate that our method achieves outstanding positioning accuracy across various walking modes and device configurations. Even when users remove headphones or a smartwatch or perform natural limb movements, the model adapts effectively, maintaining stable performance despite device removal or external interference. This highlights its strong practicality for real-life applications. Compared to our previous work Suite-IN[3], Suite-IN++ significantly reduces ATE and RTE in various walking modes, ATE and RTE are reduced by 1.02m (31.86%) and 1.33m (31.31%) for PVW, 1.43 m (33.3%) and 1.22 m (25.10%) for MVW. These results highlight the importance of decoupling global and local motion information and leveraging their complementary contributions to improve positioning accuracy. Furthermore, our innovative approach to combining global and local motion features provides a new paradigm for motion analysis in multi-device fusion, offering significant potential for future advancements in wearable-based localization systems.

### v2

@@ IV-C Weighted global fusion for overall motion trends capture
Inspired by[xue2019deepfusion], our model adopts a weighted fusion method to estimate the information quality (quality weight) contributed by each device and aggregate the motion features from all wearable devices in a weighted combination manner. Through this operation, our model aims to aggregate the effective global motion information contained in devices from different body parts into a shared low-dimensional latent space, so that the overall motion trend can be more robustly estimated without being affected by flexible and changeable device configurations.
@@ IV-D Attentive local analysis for motion details acquisition
| IONet*[chen2018ionet] | RoNIN*[herath2020ronin] | IMUNet*[zeinali2024imunet] | DeepIT*[gong2021robust] | ReWF[10447298] | Deep Sense[yao2017deepsense] | Suite-IN[sun2024suite] | Suite-IN++ | ||
@@ V Experiment
In this section, we evaluate the proposed method and compare its performance with state-of-the-art techniques. Unless explicitly stated otherwise, all experiments share the same dataset configuration, using data exclusively from user1. The dataset is divided into training, validation, and test sets in a 6:2:2 ratio, with each set covering all walking modes included in the experiments. In the multi-sensor fusion effectiveness evaluation described in Section V-C, the test set is expanded to include two additional walking modes, PK and BG, which do not appear in the training data. This setting is designed to assess the model’s ability to generalize to unseen motion modes. In the cross-subject evaluation presented in Section V-E, the test set consists of data from four users other than user1, enabling us to evaluate the model’s generalization capability across different individuals.
@@ V-A Experimental Setup
The architecture was implemented with PyTorch and trained on a NVIDIA NTX 4070 GPU. We used Adam, a first-order gradient-based optimizer[kingma2014adam], with a learning rate of 0.0001, a batch size of 128, and a window size of 100. On average, the training converged after 100 iterations. To avoid overfitting, we collected data with rich motion features and adopted Dropout [srivastava2014dropout]in the network, randomly dropping 20% of the units from the neural network during training. We set the hyper-parameters, , , , , , , .
We employ three standard metrics, as proposed in[zhang2018tutorial], to rigorously evaluate our results:
@@ V-B Compared Algorithms
Pedestrian Dead Reckoning (PDR)[tian2015enhanced]: We utilize a step-counting algorithm to detect foot-steps and move the position along the device heading direction by a predefined distance of 0.67m per step.
@@
IONet*[chen2018ionet]: A deep learning-based inertial navigation method employing an LSTM network model. To extend IONet for three-node positioning, we concatenate the data from three devices, referring to this variant as IONet*. Notably, the original IONet regresses distance and heading changes within each window; we adapt this to estimate velocity, ensuring consistency with our method.
RoNIN*[herath2020ronin]: A deep learning-based inertial tracking method that employs three different backbones (LSTM, ResNet, TCN), with ResNet achieving the highest accuracy. We extend the official RoNIN (ResNet) implementation by applying a concatenation approach for three-node data fusion, denoted as RoNIN*.
IMUNet*[zeinali2024imunet]: IMUNet introduces a one-dimensional version of the state-of-the-art convolutional neural network (CNN) network for inertial position estimation on the edge device implementation. We extend it to three-node positioning through concatenation, named as IMUNet*.
@@
DeepIT*[gong2021robust]: An inertial navigation method that integrates smartphone and headphone data using an LSTM. We extend it with DeepIT* to fuse data from three sensors with primal weighting and modify the regression target from distance and heading to velocity for consistency with our method.
@@
DeepSense*[yao2017deepsense]: DeepSense is the classic learning model for HAR of multi-sensor data. The architecture of DeepSense includes three layers of local CNN, three layers of global CNN and two layers of GRU. We implement DeepSense* for positioning based on the settings in the original article.
Suite-IN[sun2024suite]: Our previous work, a three-node positioning method based on wearable devices that uses shared global motion information contained in multiple devices.
@@ V-C Comparison with the State-of-the-Art methods
Specifically, in the stable walking mode STW, our method achieves comparable positioning results with the classic single-node localization algorithm IMUNet*[zeinali2024imunet], and outperforms multiple other localization methods. As our algorithm is specially designed to improve the positioning accuracy for complex walking modes, our method shows significant advantages in challenging scenarios (PVW, MVW, DLW, and DRW). Suite-IN[sun2024suite]achieves suboptimal performance in overall experiment. Compared with Suite-IN, our method significantly reduces ATE and RTE in all walking modes, ATE and RTE are reduced by 0.12m (3.83%) and 0.44m (11.03%) for STW, 1.02m (31.86%) and 1.33m (31.31%) for PVW, 1.43 m (33.3%) and 1.22 m (25.10%) for MVW, 0.07 m (2.12%) and 0.31 m (8.89%) for DLW, 0.11 m (4.21%) and 0.59 m (15.94%) for DRW. Compared with DeepSense[yao2017deepsense], a classic multi-sensor fusion method for human motion analysis, our approach achieves significant improvements even in complex walking modes. For MVW, DLW, and DRW, ATE is reduced by 3.15m, 2.00m, and 3.05m, while RTE decreases by 3.01m, 1.87m, and 3.18m, respectively. These results highlight the effectiveness of our method in localization under complex motion conditions.
@@
| No. | Modules | Data setting | ||||||||||||||||
| Contrast.FE | Weighted.GF | Attentive.LA | Overall | STW | PVW | MVW | DLW | DRW | ||||||||||
| ATE | RTE | ATE | RTE | ATE | RTE | ATE | RTE | ATE | RTE | ATE | RTE | |||||||
| (1) | 3.229 | 3.627 | 3.279 | 3.435 | 2.397 | 3.548 | 3.717 | 4.133 | 3.689 | 3.828 | 2.436 | 3.184 | ||||||
3.016 | 3.524 | 3.068 | 3.543 | 2.380 | 3.485 | 3.573 | 4.228 | 3.161 | 3.363 | 2.455 | 2.925 | |||||
3.209 | 3.642 | 3.129 | 3.392 | 2.095 | 3.183 | 3.367 | 4.074 | 3.635 | 3.499 | 2.863 | 3.744 | ||||
3.383 | 3.740 | 3.352 | 3.633 | 2.387 | 3.222 | 4.113 | 4.124 | 3.569 | 3.477 | 2.978 | 4.085 | ||||
4.101 | 4.194 | 3.749 | 3.952 | 3.461 | 4.329 | 5.029 | 4.935 | 5.006 | 4.101 | 3.169 | 3.836 | ||||
2.915 | 3.253 | 3.062 | 3.542 | 2.188 | 2.909 | 2.865 | 3.628 | 3.147 | 3.156 | 2.434 | 3.083 | |||
@@ V-D Ablation study
Based on variant 3), 6) introduces the Weighted.GF module to enhance global motion feature utilization. The results show that this module further improves the positioning accuracy in various walking modes, particularly in complex scenarios, as global features offer a more stable motion pattern, mitigating the negative impact of limb shaking on local features. Combining global and local motion information enables a more comprehensive capture of walking characteristics, enhancing robustness in challenging environments.Fig.9 illustrates the sensitivity of the Weighted.GF module to the hyperparameters in Equation11. When and are around the order of , the model achieves an optimal balance between global trajectory consistency and local stability. When is approximately , it effectively suppresses local pose drift and improves short-term relative positioning accuracy. Beyond these ranges, the model’s performance tends to stabilize, indicating clear parameter-sensitive regions. This parameterized design enables the model to dynamically adjust sensor weights according to different walking scenarios, achieving robust localization.
@@ V-E Tests Involving Multiple Subjects
We selected the single-node positioning method IMUNet*[zeinali2024imunet] and the multi-node positioning method Suite-IN[sun2024suite] as benchmark comparisons to evaluate the performance advantages of our method, as they perform outstandingly in their respective categories according to Tab. III. As shown in Fig. 10, our method can still maintain high positioning accuracy on unseen users without significant error growth, while the positioning performance of IMUNet and Suite-IN on unseen users is significantly reduced, indicating that they are less adaptable to the motion patterns of different individuals. In contrast, our method can more comprehensively model and adapt to the motion patterns of different individuals by effectively decoupling and utilizing different motion features, thereby significantly improving the generalization ability of the model.
@@ V-F Comparison of Suite-IN++ and Suite-IN
In this section, we compare two wearable-based pedestrian localization methods: Suite-IN++ and Suite-IN[sun2024suite]. Unlike our previous work, Suite-IN, which treats local motion features as interference, Suite-IN++ leverages the rich motion information within these features to enhance position estimation.
@@
| Method | Inference speed (fps) | FLOPs (G) | Parameters (K) |
| IONet | 44441 | 0.004 | 114.6 |
| IONet* | 46036 | 0.004 | 119.2 |
| RoNIN | 9906 | 0.021 | 4438.2 |
| RoNIN* | 10337 | 0.021 | 4443.7 |
| IMUNet | 9462 | 0.020 | 4068.2 |
| IMUNet* | 9682 | 0.020 | 4073.6 |
| DeepIT* | 20603 | 0.005 | 317.3 |
| ReWF | 5902 | 0.223 | 14049.1 |
| DeepSense | 17379 | 0.007 | 318.2 |
| Suite-IN | 8008 | 0.026 | 104.3 |
| Suite-IN++ | 9381 | 0.020 | 498.8 |
IONet, RoNIN, and IMUNet are originally designed based on smartphone data, whereas IONet*, RoNIN*, and IMUNet* are extended to support pedestrian localization using data from three wearable devices.
V-G Model Efficiency and Deployment Analysis
Although Suite-IN++ incorporates multiple CNNs, GRUs, and attention mechanisms, it was intentionally designed to support mobile deployment. To comprehensively evaluate the model’s deployability, we selected three complementary indicators: (1) inference speed (fps), which reflects real-world execution efficiency; (2) FLOPs, which represent the theoretical computational complexity; and (3) the number of parameters, indicating memory and storage requirements. Together, these metrics provide a balanced and multi-dimensional assessment of the model’s suitability for mobile computing environments.
As shown in Tab. VI (Table VI in the revised manuscript), Suite-IN++ requires only 0.020 GFLOPs and 498.8K parameters, and achieves an inference speed of 9,381 fps on a desktop GPU. These results indicate low computational overhead and excellent real-time performance, making Suite-IN++ suitable for deployment on resource-constrained devices such as smartphones. This efficiency is achieved through deliberate architectural optimization: lightweight CNNs with pooling compress temporal features, GRUs and attention modules operate on low-dimensional embeddings, and the overall network maintains moderate depth to balance complexity and modeling capacity.
In contrast, RoNIN*[herath2020ronin] and ReWF[10447298] exhibit similar FLOPs to Suite-IN++, but their parameter counts are significantly higher—4,443.7K and 14,049.1K respectively. This results in larger model sizes and higher memory usage, which may hinder their deployment on mobile devices. Meanwhile, models like DeepSense[yao2017deepsense] and DeepIT*[gong2021robust] are more lightweight, but their early fusion design and shallow temporal modeling limit their ability to handle multi-sensor inputs and complex walking patterns effectively.
Moreover, compared with our previously proposed Suite-IN[sun2024suite], Suite-IN++ significantly reduces computational complexity and improves inference speed through architectural optimization, while maintaining a moderate number of parameters. By further enhancing localization accuracy, it achieves a better balance between performance and efficiency, thereby improving its practicality for real-time deployment on mobile devices.
@@ VI CONCLUSIONS
This paper introduces a flexiwear bodynet-based inertial dataset covering flexible device configurations and complex walking modes, and proposes an innovative inertial positioning method based on real-life wearable devices. By extracting and integrating both global and local motion features, our method effectively captures the overall motion trend and detailed dynamics, leveraging sensor data from various parts of the body to achieve robust and high-precision pedestrian positioning. Experimental results demonstrate that our method achieves outstanding positioning accuracy across various walking modes and device configurations. Even when users remove headphones or a smartwatch or perform natural limb movements, the model adapts effectively, maintaining stable performance despite device removal or external interference. This highlights its strong practicality for real-life applications. Compared to our previous work Suite-IN[sun2024suite], Suite-IN++ significantly reduces ATE and RTE in various walking modes, ATE and RTE are reduced by 1.02m (31.86%) and 1.33m (31.31%) for PVW, 1.43 m (33.3%) and 1.22 m (25.10%) for MVW. These results highlight the importance of decoupling global and local motion information and leveraging their complementary contributions to improve positioning accuracy. Furthermore, our innovative approach to combining global and local motion features provides a new paradigm for motion analysis in multi-device fusion, offering significant potential for future advancements in wearable-based localization systems.
Lan Sun (Student Member. IEEE) received the B.S. degree in Measurement and Control Technology and Instrument from Southeast University, Nanjing, China, in 2023. She is currently pursuing the Ph.D. degree with Shanghai Jiao Tong University, Shanghai, China. Her current research interests include machine learning, inertial navigation, multi-sensor fusion and wearable sensor-based human motion analysis. |
Songpengcheng Xia (Member. IEEE) received the B.S. degree in navigation engineering from Wuhan University, Wuhan, China, in 2019. He is currently pursuing the Ph.D. degree with Shanghai Jiao Tong University, Shanghai, China. His current research interests include machine learning, inertial navigation, multi-sensor fusion and sensor-based human activity recognition. |
Jiarui Yang (Student Member, IEEE) received the B.S. degree in telecommunication engineering from Politecnico di Torino, Turin, Italy in 2018. He received the M.S. degree in communication systems from KTH Royal Institute of Technology, Stockholm, Sweden in 2021. He is currently working toward the Ph.D. degree in information and communication engineering with Shanghai Jiao Tong University, Shanghai, China. His research interests include machine learning, deep learning, and human pose estimation. |
Ling Pei (Senior Member, IEEE) received the Ph.D. degree from Southeast University, Nanjing, China, in 2007. From 2007 to 2013, he was a Specialist Research Scientist with the Finnish Geospatial Research Institute. He is currently a Professor at the School of Electronic Information and Electrical Engineering, Shanghai Jiao Tong University. He has authored or co-authored over 100 scientific papers. He is also an inventor of 25 patents and pending patents. His main research is in the areas of indoor/outdoor seamless positioning, ubiquitous computing, wireless positioning, Bio-inspired navigation, context-aware applications, location-based services, and navigation of unmanned systems. Dr. Pei was a recipient of the Shanghai Pujiang Talent in 2014 and ranked as the World’s Top 2% scientists by Stanford University in 2022. |
![[Uncaptioned image]](lansun.jpg)
![[Uncaptioned image]](xspc1.png)
![[Uncaptioned image]](YJR.png)
![[Uncaptioned image]](Ling_Pei.png)